Circular No. 8390

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004dj in NGC 2403 D. Pooley and W. H. G. Lewin, Massachusetts Institute of Technology,

on behalf of a larger collaboration, report the detection of x-ray emission from the type-IIP supernova 2004dj (IAUC 8377, 8378, 8379) with the Chandra X-ray Observatory: “A 40900-s ACIS-S3 observation on Aug. 9.3 UT revealed a source at α = 7^{h}37^{m}17^{s}.05, δ = +65^{o}35′57′′.9 (equinox 2000.0; ± 0′′.5 in each coordinate), in agreement with the reported optical and radio positions. Five hundred x-ray photons were detected, and preliminary spectral fits (using a MEKAL warm-plasma model) indicate a temperature of kT = 7 keV with a fitted column density of N_{H} = 6.6 × 10^{20} cm^{−2}. The unabsorbed 0.5- to 8-keV luminosity calculated from the model is 1.5 × 10^{38} erg/s (for a distance of 3.2 Mpc). This x-ray luminosity is about three times that of the type-IIP SN 1999em and nearly fifteen times that of the type-IIP SN 1999gi at similar epochs. A previous Chandra observation of NGC 2403 on 2001 Apr. 17 showed no x-ray emission near the position of SN 2004dj, with a limiting luminosity of ≈ 10^{36} erg/s. We encourage continued monitoring at all wavelengths. We plan to observe this source again with Chandra around Aug. 20, Oct. 1, and 2005 Jan. 15.”

SUPERNOVAE 2004df, 2004di, 2004dp, 2004dv, AND 2004ds A. V. Filippenko, R. J. Foley, and M. Ganeshalingam, University of

California, Berkeley, report that inspection of CCD spectra (range 330– 1000 nm), obtained on Aug. 16 UT with the Shane 3-m reflector at Lick Observatory, reveals that both SN 2004df (IAUC 8374) and SN 2004dp (IAUC 8383) are of type II (probably II-P, ∼ 2 months after the explosion, based on the well-developed P-Cyg Hα profiles). SN 2004di (IAUC 8376) is of type Ia; specifically, its spectrum closely resembles that of the peculiar, subluminous SN 1991bg (Filippenko et al. 1992, A.J. 104, 1543) ∼ 3 weeks past maximum brightness. SN 2004dv (IAUC 8388) is of type II, many months past maximum brightness; it is well into the nebular phase. SN 2004ds (IAUC 8386) is of type II; the well-developed P-Cyg hydrogen Balmer-line profiles suggest a type II-P subclassification, and an age of roughly 1 month after the explosion.

COMET 29P/SCHWASSMANN-WACHMANN Visual total-magnitude estimates by J. J. Gonzalez, Leon, Spain (0.20-

m reflector): June 30.10 UT, 11.5; July 13.10, 12.2; Aug. 13.09, 12.8.

2004 August 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green